OpenStudy (anonymous):

We say that a solution n of a congruence is unique modulo c if any solution n' of it is congruent to n modulo c. Can you give me an example?

OpenStudy (zzr0ck3r):

\(3x\equiv1(\text{mod }11)\) in \(\mathbb{Z}_{11}\)

OpenStudy (zzr0ck3r):

only solution is \(4\), but if we change to \(\mathbb{Z}\) there are infinite solutions.

OpenStudy (anonymous):

all solutions are x = 4 + 11k, right?

OpenStudy (zzr0ck3r):

in \(\mathbb{Z}\)?

OpenStudy (zzr0ck3r):

\(3(4+11k)-1=12+33k-1=11+33k=11(3k+1)\) yep